Given the μ = 24 and σ = 3, compute the z-score that corresponds to the given score X = 39


1
Expert's answer
2022-02-02T13:12:11-0500
"Z=\dfrac{X-\mu}{\sigma}=\dfrac{39-24}{3}=5"